OpenStudy (anonymous):

Write x2 - 2x - 3 = 0 in the form (x - a)2 = b, where a and b are integers.

OpenStudy (anonymous):

The answer choices are: (x - 4)2 = 3 (x - 3)2 = 2 (x - 2)2 = 1 (x - 1)2 = 4 Can someone also explain how to do it please?

OpenStudy (mathmath333):

ok

OpenStudy (anonymous):

x^2-2x-3=0 becomes (x-3)(x+1)=0, and x-3=0 Multiply 2 both sides, (x-3)2=0

OpenStudy (anonymous):

That isn't one of the answer choices. Do you mean (x-3)^2=1 ?

jhonyy9 (jhonyy9):

@Probhat prove it please that is right answer

jhonyy9 (jhonyy9):

so because not is sure

jhonyy9 (jhonyy9):

@mathmorelikeno try the last choice ,calcule it and see what you will get

OpenStudy (anonymous):

no..Check ur ans. x values can either 3 or -1 Ur choices, 1. 2x-8=3, x=11/2 2. 2x-6=2, x=8/2=4 3. x=5/2 4. x=4 Ans may be wrong!!!

jhonyy9 (jhonyy9):

@mathmorelikeno hope you know that (x-1)^2 = x^2 -2x +1 ,yes but the last term need being -3 yes ? so this mean that you need assuming to +1 minus 4 and will get -3 ok ?

OpenStudy (anonymous):

oh! last one x=3

jhonyy9 (jhonyy9):

@mathmorelikeno do you understand it now ?

OpenStudy (anonymous):

soo it would be (x - 3)^2 = 2 because in the format (x - a)2 = b, it would be (x-3)^2=2

jhonyy9 (jhonyy9):

so than you calcule (x-3)^2 =2 what you will get ??? so the needed answer sure ?

OpenStudy (anonymous):

you don't need to calculate it though

jhonyy9 (jhonyy9):

but you need to prove it

jhonyy9 (jhonyy9):

so in case of (x-3)^2 =2 c will be 9-2=7 but need being -3

OpenStudy (anonymous):

(x-a)2=b and (x-a)^2=b are not same. Clearify ur question plz.

jhonyy9 (jhonyy9):

so yes he not can writing again exponent

OpenStudy (anonymous):

@Probhat I meant (x-a)^2=b in the initial question, sorry. @jhonyy9 So it would be B?

jhonyy9 (jhonyy9):

why B ? calcule it and see what will get ,ok ?

jhonyy9 (jhonyy9):

so than you see above i have said to you try calcule the last one ,yes ?

OpenStudy (anonymous):

kk. den..last one may be d ans. calculate it. done!!

jhonyy9 (jhonyy9):

try it here for can understanding @mathmorelikeno too

jhonyy9 (jhonyy9):

please

jhonyy9 (jhonyy9):

so (x-1)^2 = x^2 -2x +1 but you know that c ,the last term need being -3 so than you need assuming to "1 -4 what is equal -3 yes ? so than you need to make equal (x-1)^2 = 4

OpenStudy (anonymous):

Answers are in to the power 2 so we have to make the given one to to the power 2. So, x^2-2x-3=0 x^2-2.x.1+1-3-1=0 (x-1)^2=4